EVIDENCE FOR A STRONG INTERMOLECULAR BOND IN THE PHENOL·N_{2} COMPLEX CATION

\underline{MARTIN C. R. COCKETT}, DARREN M. CHAPMAN, CAROLINE E. H. DESSENT, WOLF D. GEPPERT, STEPHEN R. HAINES, KLAUS MÜLLER-DETHLEFS, and MARK J. WATKINS, Department of Chemistry, University of York, Heslington, York, United Kingdom, YO1 5DD.

The vibrational spectrum of the [phenol·N_{2}]^{+} complex cation has been observed for the first time by high resolution two-colour threshold ionisation techniques (ZEKE-PFI spectroscopy and mass analysed threshold ionisation (MATI) spectroscopy). The spectra exhibit vibrational progressions in three modes, which have been assigned on the basis of ab initio calculations, to the three intermolecular modes (in-plane bend β′ 35 cm^{−1}, wag γ′ 113 cm^{−1} and stretch σ′ 125 cm^{−1}). ZEKE and MATI spectra were acquired through different intermediate S_{1} vibrational levels to aid the assignment of the vibrational structure in the cation spectra. Features were observed in the [phenol·N_{2}]^{+} MATI channel up to an ion internal energy of 1400 cm^{−1} providing a lower bound for the dissociation energy of the cationic complex. Although the binding energy is weaker than the hydrogen bonded [phenol·water] cluster^{a}, it is considerably stronger than the van der Waals [phenol·Ar]^{+} complex^{b} suggesting an intermediate binding case characteristic of a combined charge-induced dipole / charge-quadrupole interaction.

aG. Reiser, O. Dopfer R. Lindner, G. Henri, K. Müller-Dethlefs, E. W. Schlag, and S. D. Colson, Chem. Phys. Lett. \underline{181}, 1 (1991).

bE. Beiske, M. W. Rainburn, I. M. Atkinson, and A. E. W. Knight, J. Chem. Phys. \underline{91}, 752 (1989).